If the length of a rectangle is increased by 50%, by what % must the width be decreased if the area is to remain the same?
Answer by Alan3354(69443)

If the length of a rectangle is increased by 50%, by what % must the width be decreased if the area is to remain the same?
-------------------------
Increased by 50% is times 1.5, or times 3/2.
For the product L*W (= area) to remain the same, the Width must be 2/3, or decreased by 1/3.
Area = LW original
Area = (3/2)L * (2/3)W = LW